OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). 1. y - 4 = -1/2(x + 6) 2. y - 4 = 2(x + 6) 3. y + 6 = -1/2(x - 4) 4. y + 6 = 2(x - 4)

OpenStudy (anonymous):

3 is the answer, because the slope is -1/2 and then y-y1= y+6. Also x-x1= x-4. hope it helps :)

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

np anytime

OpenStudy (anonymous):

Could you help me with another question?

OpenStudy (anonymous):

sure I'll try

OpenStudy (anonymous):

Okay, here it is: Given the equation y - 3 = 1/2 (x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = 1/2 x y = 1/2 x + 6 y = 1/2 x + 9

OpenStudy (anonymous):

The only one in standard form is x-2y=-12, also I did the math and I think that is the answer.

OpenStudy (anonymous):

Okay, thanks

OpenStudy (anonymous):

Wait, it said to put it in point-slope form though

OpenStudy (anonymous):

Oh...never mind, I read it wrong. Thanks for helping me!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I got the first question wrong, but that's alright

OpenStudy (anonymous):

oh so sorry I got x and y mixed up

OpenStudy (anonymous):

it should have been 2 sorry

OpenStudy (anonymous):

That's okay, you still helped me with the second one :)